OpenStudy (anonymous):

Hertz charges 25 dollars for each mile and enterprise charges 50 dollars flat rate and 10 for every mile , at what mile would you pay the same price for enterprise and hertz

OpenStudy (anonymous):

25m=10m+50 15m=50 m=50/15=3 1/3 miles